OpenStudy (anonymous):

Write the function y-3=2/3(x-4) in the equivalent form most appropriate for identifying the slope and y-intercept of the function.

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@israa88

OpenStudy (anonymous):

what grade is this for

OpenStudy (anonymous):

9th same class as me

OpenStudy (anonymous):

oh yeah lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@urbanmorgans

OpenStudy (anonymous):

@Australopithecus

OpenStudy (anonymous):

honestly i am not sure so i dont want to give u a wrong answer

OpenStudy (anonymous):

It's fine. Thank you anyways @israa88

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

ok i think you need to do it in y=mx+b

OpenStudy (anonymous):

@esam2

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@edwardp

OpenStudy (anonymous):

\[y-3=\frac{2}{3}(x-4) \] right?

OpenStudy (anonymous):

Yes, sorry. @satellite73

OpenStudy (anonymous):

 wanna find y-intercept, plugin x=0 and solve for y